OpenStudy (anonymous):

Integral of sqrt(sin^2(x)cos^2(x))dx from 0 to 2pi

OpenStudy (freckles):

\[\sqrt{[\sin(x)\cos(x)]^2}=\sin(x)\cos(x) \text{ if } \sin(x)\cos(x)>0 \\ \sqrt{[\sin(x)\cos(x)]^2}=-\sin(x)\cos(x) if \sin(x)\cos(x)<0\]

OpenStudy (freckles):

so looking at the unit circle between 0 and 2pi both sin and cos are + in the first quadrant so the product of sin and cos is positive they are opp in the second quadrant so the product will be neg they are same signed in third quadrant so their product will be pos they are opp in the fourth so the product will be neg

OpenStudy (freckles):

so what I'm trying to say is this: \[\int\limits_{0}^{\frac{\pi}{2}}\sin(x)\cos(x) dx -\int\limits_{\frac{\pi}{2}}^{\pi}\sin(x)\cos(x) dx\\+\int\limits_{\pi}^{\frac{3\pi}{2}}\sin(x)\cos(x)dx-\int\limits_{\frac{3\pi}{2}}^{2\pi} \sin(x)\cos(x) dx\]

OpenStudy (anonymous):

oh i see

OpenStudy (freckles):

I used the definition of absolute value since |x|=sqrt(x^2)

OpenStudy (freckles):

now sin(x)cos(x) is an odd function since sin(-x)cos(-x)=-sin(x)cos(x) So that means we can probably write this as one integral times 4

OpenStudy (freckles):

\[4 \int\limits_{0}^{\frac{\pi}{2}}\sin(x)\cos(x) dx \]

OpenStudy (freckles):

either way though

OpenStudy (anonymous):

oh, realizing the odd thing made it a lot simpler, i already did all 4 and combined them :P

OpenStudy (freckles):

lol it is fine

OpenStudy (anonymous):

Either way, it's 2, which is what my teacher and calculator said :)